OpenStudy (anonymous):

find the slope of the line through the pair of point (-1/4,1/3) and (-1/2,0) fractions confuse me help please

OpenStudy (austinl):

Ok, so do you recall the formula that I gave you?

OpenStudy (anonymous):

yes but with fractions how would i simplify \[\frac{0+(-1/2) }{1/3+(-1/4) }=\]

OpenStudy (anonymous):

to find slope (m) , let point be (x1,y1) and (x2,y2) then slope, m =y2-y1/x2-x1 or y1-y2/x1-x2 m=(0-(1/3))/((-1/2)-(-1/4))

OpenStudy (anonymous):

you are not suppsed to add the coordinates .. you have to subtract them.

OpenStudy (austinl):

\[m=\frac{y_2-y_1}{x_2-x_1}\] So, I get \[m=\frac{\frac{1}{3}-0}{-\frac{1}{4}-(-\frac{1}{2})}\]

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

i dont think whhat i got is right

OpenStudy (austinl):

Ok, so the next part is the fraction math. What is \[\frac{1}{3}-0\] and \[-\frac{1}{4}-(-\frac{1}{2})\]

OpenStudy (anonymous):

answer is 4/3

OpenStudy (austinl):

You shouldn't just give answers, you should show them so that they can do it again on their own.

OpenStudy (anonymous):

so 1/3-0 would equal to 3 then -1/4-(-1/2)=1/4 and then it would be 4

OpenStudy (austinl):

1/3-0=1/3 the second half is correct at 1/4 so, we now have \[\frac{\frac{1}{3}}{\frac{1}{4}}\] how do we solve that?

OpenStudy (anonymous):

numberator we have=1/3. your denominator is -1/4-(-1/2) = -1/4+1/2 we an write that as (-1+2)/4=1/4 that is nuthng but 1/3*4

OpenStudy (austinl):

For the final part we would do this \[\frac{1}{3}\times \frac{4}{1}\]

OpenStudy (anonymous):

then we would get 4/3 now i see thank you so much are u a teacher?

OpenStudy (austinl):

Fractal division is merely multiplying by the reciprocal of the denominator. And no I'm not a teacher, I am a freshman college student.

OpenStudy (anonymous):

ur good a math i such

OpenStudy (anonymous):

suck

OpenStudy (austinl):

It just takes practice, and lots of it. If you do these problems enough it becomes second nature. Just keep up the good work.

OpenStudy (anonymous):

ok thank you